satellite73 (satellite73):

what snappy reasoning shows that \[3^{100}-2^{100}\] is divisible by 11.

OpenStudy (anonymous):

i found a solution by factoring but it was not at all obvious, so i hope there is a simple 'splanation

OpenStudy (anonymous):

U mean apart from your amazing feat of mental calculation just recently:-)

OpenStudy (anonymous):

well i factored. what mistake did i make now?

OpenStudy (anonymous):

Just kidding...

OpenStudy (anonymous):

i factored \[x^{100}-y^{100}\] and one of them, when you make the replacements, gives 55. but i thought there might be some simpler snappier number theoretic ish reason

OpenStudy (anonymous):

Doing one at a time is no problem, I'm just mulling over the subtraction.

OpenStudy (anonymous):

maybe there is a snappy way to factor. you get \[(x-y)(x+y)( \text{ this one gives 55})(\text{other stuff})\]

OpenStudy (anonymous):

Maybe u can do it: 3^100 = (3^10)^10 = 1 mod 11 and 2^100 = (2^10)^10 = 1 mod 11 by FLT and I see no reason not to subtract these in terms of the remainders, 1-1 = 0 mod 11

OpenStudy (anonymous):

There is only one factor divisible by 11. Refer to the attached Mathematica calculation.

OpenStudy (anonymous):

The number theory method isn't concerned with the factors only the remainders.

OpenStudy (anonymous):

thanks! i was fairly certain there was a snap way to do this. thanks!